OpenStudy (anonymous):

0.01 = 1–1 1–2 10–1 10–2

OpenStudy (anonymous):

Rewrite it back words without the decimal.

OpenStudy (anonymous):

Then multiply them like this But It can't be the first two because they don't have the base of 10.

OpenStudy (anonymous):

So the last two are 10*1=10 10*10=100

OpenStudy (anonymous):

So if you write 001 backward what would it be?

OpenStudy (anonymous):

100 lol

OpenStudy (anonymous):

yep :) also a number multiplied by 10 can't be greater or less then 1-10..just though I should say.

OpenStudy (anonymous):

I meant the base of 10

OpenStudy (anonymous):

0.01 = 1/100 = 1/10^2 = 10^(-2)